# Atherosclerosis through Hierarchical Explainable Neural Network Analysis

### v1

@@ 4.2 Classification and Profiling Performance
Sanchez-Cabo et al. [27] Steenman et al. [30]
@@
Deep Learning Approaches
MLP (3-layer) 0.712 ± 0.05 0.533 ± 0.06 0.805 ± 0.007 0.786 ± 0.012
MLP (5 layer) 0.708 ± 0.03 0.525 ± 0.03 0.827 ± 0.010 0.808 ± 0.009
Graph Deep Learning (Patient Similarity)
@@
Hierarchical Network Classification (ATHENA)
GCN 0.812 ± 0.062 0.725 ± 0.072 0.837 ± 0.013 0.832 ± 0.011
GAT 0.808 ± 0.058 0.718 ± 0.070 0.843 ± 0.009 0.839 ± 0.010
GraphSage 0.805 ± 0.060 0.721 ± 0.068 0.839 ± 0.011 0.836 ± 0.014
Table 2: Performance Metrics of Omics-based Patient Classification Models. Comparison of classification performance across multiple models and feature combinations using metrics tested. Results are presented as mean ± standard deviation over 5 fold stratified cross validation. Models incorporating the Hierarchical Network representation (ATHENA) to generate graph topology achieved the highest performance across all metrics.
In Table 2, we examined the performance of our hierarchical network representation (ATHENA) against linear modeling (Lasso, Elastic-Net with SAGA optimizer and an L1 ratio of 0.5, and Support Vector Classifier (SVC) [10]), deep learning (3-layer and 5-layer multi-layer perceptron (MLP) [10]), and graph deep learning (Graph Convolutional Network [11], Graph Attention Network [39], and GraphSage [7]) methods. The mean and standard deviation of AUC and F1 score for all methods are summarized (Table 2), with the best performing models highlighted in green. In general, graph-based approaches outperform non-graph methods, and integrating molecular interaction topology alongside clinical features leads to added performance gains. Notably, ATHENA’s hierarchical network representation contributed to the uniform performance gains across all graph deep learning methods evaluated.
@@
Figure 3: Model performance against individual subtypes of atherosclerosis for Sanchez-Cabo et al. [27]. Upper panels show the consistent performance of ATHENA across all 3 subtypes in their mean AUC curves. Lower panels show the t-SNE-based separation of 2 distinct clusters in each subtype of atherosclerosis. Each dot represents one patient. For generalized atherosclerosis, cluster 1 has 27 patients and cluster 2 has 22 patients. For intermediate atherosclerosis, cluster 1 has 34 patients and cluster 2 has 27 patients. For focal atherosclerosis, cluster 1 has 75 patients and cluster 2 has 71 patients.
@@
Sanchez-Cabo et al. [27] Steenman et al. [30]
@@
Table 3: Ablation Results for ATHENA. Various ablation configurations of ATHENA on various classification metrics. The configurations include the Patient Similarity Network alone, the Protein-Protein Interaction (PPI) Network alone, and ATHENA. Integrating hierarchical structures and clinical features significantly improves performance, as demonstrated by ATHENA, achieving the highest scores across all metrics.
4.3 Ablation Studies
To demonstrate the effectiveness of ATHENA, we scrutinized three model configurations: (1) patient similarity network only with transcriptomic markers and clinical features, (2) PPI network only with transcriptomics markers, and (3) ATHENA. Table 3 summarizes the mean and standard deviation of AUC and F1 score for each configuration. Using only a patient similarity network representation resulted in the lowest performance, whereas relying solely on the PPI network representation produced slightly better results. The integration of both types of networks into a unified graph data representation achieved the highest overall performance across all metrics, underscoring the notion that a hierarchical network integration is the optimal method. Compared to patient similarity networks alone, ATHENA has had average AUC improvements of 0.057 to 0.065 and F1 score improvements of 0.041 to 0.042 for the cohort from Sanchez-Cabo et al. [27]. For the cohort from Steenman et al. [30], ATHENA has had average AUC improvements of 0.011 to 0.025 and F1 score improvements of 0.022 to 0.037. These results highlight that it is the synergistic representation provided by combining the two networks, not just the deep learning approach, that drives performance gains.
4.4 Discovery of Pathogenesis Clusters
@@
In the generalized subtype, two distinct clusters capture the balance between endothelial inflammation and structural remodeling. Cluster 1 is characterized by dysfunctional and inflammatory pathways, which is highlighted by up-regulation of CDK5R1 and ICAM1, markers known to contribute to endothelial regulation in atherosclerotic patients [3, 12]. In contrast, Cluster 2 is characterized by structural and remodeling pathways, exhibiting higher TTLL5 which drives microtubule remodeling necessary for thickening of the arterial wall [5]. Cluster 2 also exhibits an upregulation RASGRP3 expression, thereby favoring cytoskeletal adaptations over inflammation [33].
@@
Stories reviewed in the intermediate subtype show that Cluster 1 amplifies hallmark atherogenic processes within monocyte and inflammatory pathways with elevated ITGAM/CD11b expression, leading to thickening of the arterial wall as shown in previous atherosclerotic studies [15]. Furthermore, MMP9 upregulation is seen with CD14 enhancing LPS/TLR4 signaling in infiltrating monocytes [44, 1]. Metabolic sensors INSR and FOSL2 are also evident [29, 45]. In parallel, Cluster 2 exhibits regulatory and metabolic pathways, characterized by SMIM24, MAF, and FCER1A representing a subset of intermediate lesions engaging lipid-handling or anti-inflammatory programs to temper unchecked matrix remodeling [36, 16].
@@
In the subtype of advanced focal atherosclerosis, Cluster 1 hosts oxidative-stress and inflammatory regulation [6], pro-inflammatory cytokine release [2], and HSPB1 chaperones proteins under oxidative stress. Cluster 2 hosts fibrotic-matrix remodeling pathways, which drive fibrous-cap integrity [38], ITGB7-dependent integrin adhesion, and CEMIP2-directed hyaluronan reorganization in the extracellular matrix [47], collectively stabilizing advanced plaques.
@@
We further investigate molecular heterogeneity in the cohort from Steenman et al. [30]. Panel II of Figure 4 presents clear separation between patients of the same clinical phenotype in two distinct clusters. Cluster 1 is characterized with upregulation of HOXA6 and HOXA5, which are involved in developmental patterning and vascular integrity [9, 48]. In contrast, Cluster 2 exhibits elevated expression of SLC2A1 and SMS, genes associated with metabolic activation and cellular transport processes [40, 13]. These two clusters reveal molecular phenotypes that differ from those observed in the Sanchez et al. cohort [27], with distinctions in this cohort of one molecular phenotype primarily driven by regulation of structural proteins and a second one enriched in metabolic pathways.
Figure 4: Model performance against cohort of atherosclerosis from Steenman et al. [30]. Panel I (top) shows consistent performance of ATHENA across the three architectures: GCN, GAT, and GraphSAGE. Panel II (bottom) shows the t-SNE-based separation of 2 distinct clusters in the cohort. Each dot represents 1 patient; cluster 1 has 34 patients; and cluster 2 has 36 patients.
5 Discussion
We present a novel hierarchical network representation that integrates both molecular interaction topology and clinical features. Linear modeling approaches and standard deep learning (multi-layer perceptron) treat molecular profiles as unstructured vectors; unfortunately, they overlook the mechanistic insights of pathogenesis embedded in molecular interaction networks and pathway topologies. In contrast, graph deep learning identifies and integrates these molecular interactions toward a specific phenotype. ATHENA extends this notion by combining patient clinical features and molecular data into a unified hierarchical graph, thereby creating a more comprehensive representation of disease phenotypes.
@@ 5 Discussion
  3. 3.
    Our analysis captures molecular drivers at a single time point, limiting our assessment of how these hierarchical network patterns evolve during disease progression. Longitudinal studies tracking molecular and clinical changes over time would provide valuable insights into the dynamic nature of atherosclerosis development and enable precision in the prediction of disease trajectories.
In summary, the superior performance of the hierarchical representation compared to isolated molecular or clinical networks underscores the fundamental interconnection between systemic factors and molecular mechanisms in atherosclerosis pathogenesis. By capturing both cohort-wide patterns through patient similarity networks and individual molecular perturbations through PPI networks, our approach reflects the multiscale nature of cardiovascular diseases where transcriptomics, clinical features, and molecular dysregulation converge to drive pathology.
Importantly, ATHENA identifies molecularly informed clusters within clinically defined subtypes of atherosclerosis. Understanding these two clusters in each subtype underscore the complexity of disease pathology, indicating that patients grouped under a single clinical diagnosis may differ substantially at the molecular level. Establishing how these molecular phenotypes can meaningfully guide patient-specific treatment decisions and elucidate underlying disease mechanisms will necessitate careful clinical study design, targeted investigations, and subsequent validation in larger cohorts. Nevertheless, this novel stratification represents a critical step toward precision medicine approaches in atherosclerosis, potentially enabling earlier intervention and improved patient outcomes.

### v2

@@ 4.2. Classification and Profiling Performance
Sanchez-Cabo et al. (Sánchez-Cabo et al., 2023) Steenman et al. (Steenman et al., 2018)
@@
[…]
@@
[…]
Table 2. Performance Metrics of Omics-based Patient Classification Models. Comparison of classification performance across multiple models and feature combinations using metrics tested. Results are presented as mean ± standard deviation over 5 fold stratified cross validation. Models incorporating the Hierarchical Network representation (ATHENA) to generate graph topology achieved the highest performance across all metrics.
In Table 2, we examined the performance of our hierarchical network representation (ATHENA) against linear modeling (Lasso, Elastic-Net with SAGA optimizer and an L1 ratio of 0.5, and Support Vector Classifier (SVC) (Khan et al., 2024)), deep learning (3-layer and 5-layer multi-layer perceptron (MLP) (Khan et al., 2024)), and graph deep learning (Graph Convolutional Network (Kipf and Welling, 2016), Graph Attention Network (Veličković et al., 2017), and GraphSage (Hamilton et al., 2017)) methods. The mean and standard deviation of AUC and F1 score for all methods are summarized (Table 2), with the best performing models highlighted in green. In general, graph-based approaches outperform non-graph methods, and integrating molecular interaction topology alongside clinical features leads to added performance gains. Notably, ATHENA’s hierarchical network representation contributed to the uniform performance gains across all graph deep learning methods evaluated.
@@
Figure 3. Model performance against individual subtypes of atherosclerosis for Sanchez-Cabo et al. (Sánchez-Cabo et al., 2023). Upper panels show the consistent performance of ATHENA across all 3 subtypes in their mean AUC curves. Lower panels show the t-SNE-based separation of 2 distinct clusters in each subtype of atherosclerosis. Each dot represents one patient. For generalized atherosclerosis, cluster 1 has 27 patients and cluster 2 has 22 patients. For intermediate atherosclerosis, cluster 1 has 34 patients and cluster 2 has 27 patients. For focal atherosclerosis, cluster 1 has 75 patients and cluster 2 has 71 patients.
4.3. Ablation Studies
To demonstrate the effectiveness of ATHENA, we scrutinized three model configurations: (1) patient similarity network only with transcriptomic markers and clinical features, (2) PPI network only with transcriptomics markers, and (3) ATHENA. Table 3 summarizes the mean and standard deviation of AUC and F1 score for each configuration. Using only a patient similarity network representation resulted in the lowest performance, whereas relying solely on the PPI network representation produced slightly better results. The integration of both types of networks into a unified graph data representation achieved the highest overall performance across all metrics, underscoring the notion that a hierarchical network integration is the optimal method. Compared to patient similarity networks alone, ATHENA has had average AUC improvements of 0.057 to 0.065 and F1 score improvements of 0.041 to 0.042 for the cohort from Sanchez-Cabo et al. (Sánchez-Cabo et al., 2023). For the cohort from Steenman et al. (Steenman et al., 2018), ATHENA has had average AUC improvements of 0.011 to 0.025 and F1 score improvements of 0.022 to 0.037. These results highlight that it is the synergistic representation provided by combining the two networks, not just the deep learning approach, that drives performance gains.
4.4. Discovery of Pathogenesis Clusters
@@
Sanchez-Cabo et al. (Sánchez-Cabo et al., 2023) Steenman et al. (Steenman et al., 2018)
@@
Table 3. Ablation Results for ATHENA. Various ablation configurations of ATHENA on various classification metrics. The configurations include the Patient Similarity Network alone, the Protein-Protein Interaction (PPI) Network alone, and ATHENA. Integrating hierarchical structures and clinical features significantly improves performance, as demonstrated by ATHENA, achieving the highest scores across all metrics.
Our molecular clustering analysis shows that patients within the same imaging-defined subtype (None, Generalized, Intermediate, Focal) exhibit distinct molecular signatures, revealing opportunities for more granular patient stratification beyond categorization of anatomical features derived from imaging. Specifically, ATHENA identified two distinct molecular clusters within each imaging-defined subtype (Figure 3, Panel II), highlighting that patients with identical imaging presentations may be driven by fundamentally different underlying mechanisms. Rather than discrete imaging categories, our molecular analysis suggests atherosclerosis exists on a continuum.
@@
In the generalized subtype, two distinct clusters capture the balance between endothelial inflammation and structural remodeling. Cluster 1 is characterized by dysfunctional and inflammatory pathways, which is highlighted by up-regulation of CDK5R1 and ICAM1, markers known to contribute to endothelial regulation in atherosclerotic patients (Bai et al., 2012; Kitagawa et al., 2002). In contrast, Cluster 2 is characterized by structural and remodeling pathways, exhibiting higher TTLL5 which drives microtubule remodeling necessary for thickening of the arterial wall (Bompard et al., 2018). Cluster 2 also exhibits an upregulation RASGRP3 expression, thereby favoring cytoskeletal adaptations over inflammation (Tang et al., 2014).
@@
Stories reviewed in the intermediate subtype show that Cluster 1 amplifies hallmark atherogenic processes within monocyte and inflammatory pathways with elevated ITGAM/CD11b expression, leading to thickening of the arterial wall as shown in previous atherosclerotic studies (Lou et al., 2024). Furthermore, MMP9 upregulation is seen with CD14 enhancing LPS/TLR4 signaling in infiltrating monocytes (Yamamoto et al., 2018; Aggarwal et al., 2023). Metabolic sensors INSR and FOSL2 are also evident (Semenkovich et al., 2006; Yin et al., 2017). In parallel, Cluster 2 exhibits regulatory and metabolic pathways, characterized by SMIM24, MAF, and FCER1A representing a subset of intermediate lesions engaging lipid-handling or anti-inflammatory programs to temper unchecked matrix remodeling (Tudzynski et al., 1980; Luo et al., 2014).
@@
In the subtype of advanced focal atherosclerosis, Cluster 1 hosts oxidative-stress and inflammatory regulation (Guo et al., 2012), pro-inflammatory cytokine release (Amini et al., 2004), and HSPB1 chaperones proteins under oxidative stress. Cluster 2 hosts fibrotic-matrix remodeling pathways, which drive fibrous-cap integrity (Uyama et al., 2017), ITGB7-dependent integrin adhesion, and CEMIP2-directed hyaluronan reorganization in the extracellular matrix (Yoshida et al., 2025), collectively stabilizing advanced plaques.
@@
We further investigate molecular heterogeneity in the cohort from Steenman et al. (Steenman et al., 2018). Panel II of Figure 4 presents clear separation between patients of the same clinical phenotype in two distinct clusters. Cluster 1 is characterized with upregulation of HOXA6 and HOXA5, which are involved in developmental patterning and vascular integrity (Jing et al., 2021; Zaina et al., 2014). In contrast, Cluster 2 exhibits elevated expression of SLC2A1 and SMS, genes associated with metabolic activation and cellular transport processes (Wall et al., 2018; Li et al., 2012). These two clusters reveal molecular phenotypes that differ from those observed in the Sanchez et al. cohort (Sánchez-Cabo et al., 2023), with distinctions in this cohort of one molecular phenotype primarily driven by regulation of structural proteins and a second one enriched in metabolic pathways.
Figure 4. Model performance against cohort of atherosclerosis from Steenman et al. (Steenman et al., 2018). Panel I (top) shows consistent performance of ATHENA across the three architectures: GCN, GAT, and GraphSAGE. Panel II (bottom) shows the t-SNE-based separation of 2 distinct clusters in the cohort. Each dot represents 1 patient; cluster 1 has 34 patients; and cluster 2 has 36 patients.
5. Discussion
We present ATHENA, a novel hierarchical network representation that integrates both molecular interaction topology and clinical features. Linear modeling approaches and standard deep learning (multi-layer perceptron) treat molecular profiles as unstructured vectors; unfortunately, they overlook the mechanistic insights of pathogenesis embedded in molecular interaction networks and pathway topologies. In contrast, graph deep learning identifies and integrates these molecular interactions toward a specific phenotype. ATHENA extends this notion by combining patient clinical features and molecular data into a unified hierarchical graph, thereby creating a more comprehensive representation of disease phenotypes.
@@ 5. Discussion
  3. (3)
    Our analysis captures molecular drivers for a patient cohort consisting of multiple stages of atherosclerosis (generalized, intermediate, and focal), representative of a cross-sectional evaluation of disease progression. ATHENA’s performance is consistent across all stages as well as discovers unique molecular subgraphs and biomarkers for each stage. However, longitudinal studies tracking molecular and clinical changes over time would provide valuable insights into the dynamic nature of atherosclerosis development and enable precision in the prediction of disease trajectories.
  4. (4)
    We acknowledge that comprehensive batch effect analysis would strengthen clinical applicability, where a commitment to extensive large cohort, multicenter clinical trials validations will be necessary.
In summary, the superior performance of the hierarchical representation compared to isolated molecular or clinical networks underscores the fundamental interconnection between systemic factors and molecular mechanisms in atherosclerosis pathogenesis. By capturing both cohort-wide patterns through patient similarity networks and individual molecular perturbations through PPI networks, our approach affords an improved understanding of the multiscale nature of cardiovascular diseases where transcriptomics, clinical features, and molecular dysregulation converge to drive pathology.
Importantly, ATHENA identifies molecularly informed clusters within clinically defined subtypes of atherosclerosis. Understanding these two clusters in each subtype addresses the complexity of disease pathology, indicating that patients grouped under a single clinical diagnosis may differ substantially at the molecular level. Establishing how these molecular phenotypes can meaningfully guide patient-specific treatment decisions and elucidate underlying disease mechanisms will necessitate careful clinical study design, targeted investigations, and subsequent validation in larger cohorts. Nevertheless, this novel stratification represents a critical step toward precision medicine approaches in atherosclerosis, potentially enabling earlier intervention and improved patient outcomes.
Acknowledgements.
We would like to acknowledge Baradwaj Simha Sankar for his feedback in guiding the direction of hierarchical analysis of atherosclerosis. This work was supported by National Institutes of Health (NIH) R35 HL135772 to P.P.; U34 HG012517 to P.P. and A.B.; and the TC Laubisch Endowment to P.P. at UCLA.